Gabriel is 16 years older than sam. in 4 years gabriel will be twice as old as sam. What are their ages now?
Answer by mananth(16946)

Sam x
Gabriel x + 16
4 years later
Sam x+4
Gabriel x+16+4 =x+20
Gabriel 2 times older
x+20=2(x+4)
x+20=2x+8
2x-x=20 -8
x=12
/ 1
x=12 years Sam
x+16= 28 years Gabriel